# NetMamba: Efficient Network Traffic Classification via Pre-training Unidirectional Mamba

### v1

@@ VI-A4 Stride Cutting
VI-B NetMamba Model
@@
VI-B1 Stride Embedding
Given the stride array, we initially perform a linear projection on each stride $\mathbf{s}_{i}$ to a vector with size $\mathtt{D}_{\text{enc}}$ and incorporate position embeddings $\mathbf{E}_{\text{pos}}^{\text{enc}}\in\mathbb{R}^{N_{s}\times\mathtt{D}_{\text{enc}}}$ as shown below:
$\mathbf{X}_{0}=\left[\mathbf{s}_{1}\mathbf{W};\mathbf{s}_{2}\mathbf{W};\cdots;\mathbf{s}_{N_{s}}\mathbf{W};\mathbf{x}_{\text{cls}}\right]+\mathbf{E}_{\text{pos}}^{\text{enc}}$ (4)
@@ VI-B1 Stride Embedding
VI-B2 NetMamba Block
Recently, several variants of Mamba have been proposed to accommodate domain-specific data formats and task requirements. For instance, Vim [13] incorporates bidirectional Mamba blocks for spatial-aware understanding of vision tasks, Graph-Mamba [14] introduces a graph-dependent selection mechanism for graph learning, while MiM-ISTD [28] customizes a cascading Mamba structure for extracting hierarchical visual information. We argue that the original unidirectional Mamba design, tailored for sequence modeling, is well-suited for representation learning in sequential network traffic, offering increased efficiency through the elimination of omnidirectional scans and redundant blocks.
Therefore, we design the NetMamba encoder and decoder using unidirectional Mamba blocks, as illustrated in Figure 2. The operational process of the NetMamba block forward pass is outlined in Algorithm 1. For a given input token sequence $\mathbf{X}_{t-1}$ with a batch size $\mathtt{B}$ and sequence length $\mathtt{L}$ from the $(t-1)$-th NetMamba block, we begin by normalizing it and then projecting it linearly into $\mathbf{x}$ and $\mathbf{z}$, both with a dimension size of $\mathtt{E}$. We subsequently apply causal 1-D convolution to $\mathbf{x}$, resulting in $\mathbf{x}^{\prime}$. Based on $\mathbf{x}^{\prime}$, we compute the input-dependent step size $\mathbf{\Delta}$, as well as the projection parameters $\mathbf{B}$ and $\mathbf{C}$ having a dimension size of $\mathtt{N}$. We then discretize $\overline{\mathbf{A}}$ and $\overline{\mathbf{B}}$ using $\mathbf{\Delta}$. Following this, we calculate $\mathbf{y}$ employing a hardware-aware SSM. Finally, $\mathbf{y}$ is gated by $\mathbf{z}$ and added residually to $\mathbf{X}_{t-1}$, resulting in the output token sequence $\mathbf{X}_{t}$ for the $t$-th NetMamba block.
@@ VI-B2 NetMamba Block
VI-C NetMamba Pre-training
VI-C1 Random Masking
@@ VI-C2 Masked Pre-training
$\mathbf{X}_{\text{out}}^{\text{enc}}$ $=\mathbf{Encoder}(\mathbf{X}_{0}^{\text{vis}})\in\mathbb{R}^{\mathtt{L}_{\text{vis}}\times\mathtt{D}_{\text{dec}}}$ (6)
$\mathbf{X}_{\text{in}}^{\text{dec}}$ $=\mathbf{Unshuffle}(\mathbf{Concat}(\mathbf{X}_{\text{out}}^{\text{enc}},\mathbf{X}_{\text{mask}}))+\mathbf{E}_{\text{pos}}^{\text{dec}}$
$\mathbf{X}_{\text{out}}^{\text{dec}}$ $=\mathbf{Decoder}(\mathbf{X}_{\text{in}}^{\text{dec}})$
where the $\mathbf{Unshuffle}$ operation restores the original sequence order, and $\mathbf{E}_{\text{pos}}^{\text{dec}}\in\mathbb{R}^{\mathtt{L}\times\mathtt{D}_{\text{dec}}}$ represents decoder-specific positional embeddings. Subsequently, the mean square error (MSE) loss for self-supervised reconstruction is calculated as shown below:
@@
$\mathbf{y}_{\text{rec}}$ $=\mathbf{Shuffle}(\mathbf{X}_{\text{out}}^{\text{dec}})[\mathtt{L}_{\text{vis}}+1:\mathtt{L},\ :\ ]$
@@
where $\mathbf{y}{\text{real}}$ represents the ground-truth mask tokens, and $\mathbf{y}_{\text{rec}}$ signifies the predicted ones.
VI-D NetMamba Fine-tuning
@@
$\mathbf{X}_{\text{out}}$ $=\mathbf{Encoder}(\mathbf{X}_{0})$ (8)
$\mathbf{f}$ $=\mathbf{X}_{\text{out}}[\mathtt{L},\ :]$
$\hat{\mathbf{y}}$ $=\mathbf{MLP(\mathbf{Norm}(\mathbf{f}))}$
@@ VII-A1 Datasets
To assess the effectiveness and generalization abilities of NetMamba, we conducted experiments using six publicly available real-world traffic datasets.
@@
    CrossPlatform(Android) [29]. This dataset comprises encrypted traffic generated by the top 100 Android applications from the US, China, and India. It encompasses 254 applications with a combined size of 25.59 MB.
@@
    CrossPlatform(iOS) [29]. This dataset consists of encrypted traffic generated by the top 100 iOS applications from the US, China, and India. It includes 253 applications with a total size of 23.92 MB.
@@
    CICIoT2022 [30]. This dataset contains traffic collected from a laboratory network aimed at profiling, behavioral analysis, and vulnerability testing for various IoT devices. We utilized benign traffic from the power experiment and malicious traffic, including Denial of Service (DoS) attack and brute force attacks, with an overall size of 6.1 GB.
  4. 4.
    ISCXTor2016 [31]. This dataset encompasses application traffic utilizing the Onion Router (Tor) for encrypted communications. Tor adds an additional layer of obfuscation to communication through a distributed routing network. We utilized Tor traffic data from 8 communication categories, with a total size of 11.9 GB.
  5. 5.
    ISCXVPN2016 [32]. This dataset comprises encrypted communication traffic tunneled through Virtual Private Networks (VPN). VPNs are commonly used for circumventing censorship and hiding locations through protocol obfuscation. We utilized VPN traffic data from 7 communication categories, with a total size of 2.4 GB.
  6. 6.
    USTC-TFC2016 [33]. This dataset contains encrypted traffic consisting of both malware and benign applications. It includes 10 categories of benign traffic and 10 categories of malicious traffic, with a total size of 3.8 GB.
We have noted a significant discrepancy in the number of flows across categories within a given dataset, which adversely affects the effectiveness and generalization capabilities of a traffic understanding model. To address the data distribution imbalance, we set lower and upper limits for each category within the dataset. Any traffic category with a flow amount below the lower limit is discarded, while those exceeding one are randomly sampled to meet the criteria. The statistical details of our normalized datasets are presented in Table III.
TABLE III: The Statistical Information of Normalized Datasets
Dataset Lower Upper # Flow # Category
CrossPlatform(Android) 50 2,000 17,624 181
CrossPlatform(iOS) 50 2,000 9,536 124
CICIoT2022 200 6,000 19,497 6
ISCXTor2016 10 4,000 188 8
ISCXVPN2016 500 4,000 13,832 7
USTC-TFC2016 500 2,000 40,000 20
@@ VII-A3 Implementation Details
The NetMamba architecture features an encoder composed of 4 Mamba blocks and a decoder composed of 2 Mamba blocks. More hyper-parameter details can be found in Table IV.
@@
TABLE IV: Hyper-Parameter details of NetMamba
@@ VII-B Overall Evaluation
TABLE V: Comparison Results on CrossPlatform(Android), CrossPlatform(iOS) and CICIoT2022
@@
TABLE VI: Comparison Results on ISCXTor2016, ISCXVPN2016 and USTC-TFC2016
@@
We evaluated the performance of NetMamba in categorizing traffic using six publicly available datasets. As shown in Table V and Table VI, NetMamba demonstrates superior performance compared to all baseline methods across five datasets. However, it falls slightly short in comparison to TFE-GNN on the CICIoT2022 dataset, with a marginal 0.15 percentage point difference in accuracy and f1 score. On average, NetMamba achieves accuracy levels ranging from 0.9869 to 0.9993 and f1 scores between 0.9864 and 0.9993. Notably, NetMamba maintains the fewest parameters among all deep learning methods, underscoring its efficient yet effective capabilities in traffic representation learning.
@@ VII-B1 CrossPlatform(Android)
As indicated in Table V, NetMamba demonstrates significant improvements over existing methods on the CrossPlatform(Android) dataset. Specifically, our model achieves a notable improvement of 4.83% in accuracy and 4.64% in f1 score compared to the state-of-the-art method (ET-BERT). ET-BERT focuses solely on learning traffic representations from packet payloads, overlooking the valuable content information carried by packet headers. In contrast, NetMamba effectively models both header and payload characteristics, leading to a more comprehensive analysis of traffic patterns.
@@ VII-B4 ISCXTor2016, ISCXVPN2016 & USTC-TFC2016
As depicted in Table VII, NetMamba surpasses all existing methods across the three encrypted datasets. Particularly noteworthy is the significantly unstable performance exhibited by methods other than YaTC and YaTC(OF) across different datasets. Given the similarities between NetMamba and YaTC, we contend that a robust traffic representation scheme, which incorporates both header and payload data, alongside a well-designed pre-training task, plays a crucial role in enhancing encrypted traffic analysis capabilities.
@@ VII-C Inference Efficiency Evaluation
Figure 3: The Inference Efficiency Comparison with Other Models
Figure 4: The Inference Efficiency Comparison with Batch Size 64
To assess the inference efficiency of NetMamba, we conducted comparison experiments focusing on speed and GPU memory consumption against existing deep learning methods. As shown in Figure 3(a), NetMamba achieves the fastest inference speed among all compared methods across various input batch sizes. This speed advantage is particularly notable due to the substantial model parameters and inefficient model architecture design present in models such as ET-BERT, TFE-GNN, and FS-Net, which result in inference speeds up to 40 times slower than that of NetMamba. Even when compared with models possessing similar parameter counts, NetMamba continues to outperform YaTC and its faster variant. This superiority is primarily attributed to the lower computational complexity of Mamba compared to Transformer models. Considering a token sequence $\mathbf{X}\in\mathbb{R}^{1\times\mathtt{L}\times\mathtt{D}}$ and the default setting $\mathtt{E}=2\mathtt{D},\ \mathtt{N}=16$, the computational complexities of a global self-attention in Transformer and SSM in Mamba are as follows:
$\Omega(\text{self-attention})=4\mathtt{L}\mathtt{D}^{2}+2\mathtt{L}^{2}\mathtt{D}$ (10)
@@
It is evident that self-attention exhibits quadratic complexity to the sequence length $\mathtt{L}$, whereas SSM is linear. This computational efficiency makes NetMamba more scalable compared to Transformer-based models like YaTC and ET-BERT.
In Figure 3(b), NetMamba demonstrates lower GPU memory consumption compared to most models except FS-Net and YaTC when using large batch sizes. FS-Net relies heavily on RNNs, which necessitate linear memory with respect to sequence length, allowing for reduced memory costs. However, it should be noted that FS-Net exhibits significantly slower inference speeds and poorer classification performance. In contrast, YaTC reduces memory costs by employing a model forward trick that shortens input sequence length to a fraction of its original size, resulting in a substantial reduction in memory usage. Without such an optimization, YaTC(OF) consumes up to four times more GPU memory than NetMamba. Compared to other baselines, NetMamba achieves improved memory efficiency primarily by customizing GPU operators. These operators are designed to minimize the storage of extensive intermediate states and conduct recomputation during the backward pass.
When the input batch size is set to 64, as depicted in Figure 4, NetMamba exhibits an improvement in speed, being 2.24 times faster than the best baseline, YaTC(OF). Apart from FS-Net and memory-optimized YaTC, NetMamba surpasses other methods in terms of GPU memory utilization. In summary, NetMamba achieves the highest inference speeds among all deep learning methods while maintaining comparably low memory usage.
TABLE VII: Ablation Study of NetMamba on All Datasets
@@
w/ BiMamba1 0.9852 0.9851 0.9793 0.9793 0.9974 0.9974 0.9986 0.9986 0.9935 0.9935 0.9987 0.9987
@@
    Substituted unidirectional Mamba blocks with bidirectional ones.
@@
    Replaced the 1-dimensional stride cutting with 2-dimensional patch splitting.
Figure 5: The Performance Comparison with Other Models on Few-Shot Settings

### v3

@@ V-4 Stride Cutting
Takeaway. Our traffic representation scheme effectively retains crucial information from both packet headers and payloads, while eliminating unwanted biases through techniques such as IP anonymization, byte balancing, and stride cutting. For a detailed evaluation, please refer to § VII-D.
VI Model Details
This section details the NetMamba model architecture, along with the pre-training and fine-tuning strategies.
VI-A NetMamba Architecture
@@
VI-A1 Stride Embedding
Given the stride array, we initially perform a linear projection on each stride $\mathbf{s}_{i}$ to a vector with size $\mathtt{D}_{\text{enc}}$ and incorporate position embeddings $\mathbf{E}_{\text{enc}}^{\text{pos}}\in\mathbb{R}^{N_{s}\times\mathtt{D}_{\text{enc}}}$ as shown below:
$\mathbf{X}_{0}=\left[\mathbf{s}_{1}\mathbf{W};\mathbf{s}_{2}\mathbf{W};\cdots;\mathbf{s}_{N_{s}}\mathbf{W};\mathbf{x}_{\text{cls}}\right]+\mathbf{E}_{\text{enc}}^{\text{pos}}$ (3)
@@ VI-A1 Stride Embedding
VI-A2 NetMamba Block
Recently, several variants of Mamba have been proposed to accommodate domain-specific data formats and task requirements. For instance, Vim [12] incorporates bidirectional Mamba blocks for spatial-aware understanding of vision tasks, Graph-Mamba [16] introduces a graph-dependent selection mechanism for graph learning, while MiM-ISTD [30] customizes a cascading Mamba structure for extracting hierarchical visual information. We argue that the original unidirectional Mamba design[14], tailored for sequence modeling, is well-suited for representation learning in sequential network traffic, offering increased efficiency through the elimination of omnidirectional scans and redundant blocks. We carefully test different Mamba variants, demonstrating that the selected unidirectional Mamba is more suitable for processing network traffic. Please refer to the ablation studies for more details.
Hence, we implement the NetMamba encoder and decoder using unidirectional Mamba blocks, as illustrated in Figure 2. The operational process of the NetMamba block forward pass is outlined in Algorithm 1. For a given input token sequence $\mathbf{X}_{t-1}$ with a batch size $\mathtt{B}$ and sequence length $\mathtt{L}$ from the $(t-1)$-th NetMamba block, we begin by normalizing it and then projecting it linearly into $\mathbf{x}$ and $\mathbf{z}$, both with dimension size of $\mathtt{E}$. We subsequently apply causal 1-D convolution to $\mathbf{x}$, resulting in $\mathbf{x}^{\prime}$. Based on $\mathbf{x}^{\prime}$, we compute the input-dependent step size $\mathbf{\Delta}$, as well as the projection parameters $\mathbf{B}$ and $\mathbf{C}$ having a dimension size of $\mathtt{N}$. We then discretize $\overline{\mathbf{A}}$ and $\overline{\mathbf{B}}$ using $\mathbf{\Delta}$. Following this, we calculate $\mathbf{y}$ employing a hardware-aware SSM. Finally, $\mathbf{y}$ is gated by $\mathbf{z}$ and added residually to $\mathbf{X}_{t-1}$, resulting in the output token sequence $\mathbf{X}_{t}$ for the $t$-th NetMamba block.
@@ VI-A2 NetMamba Block
VI-B NetMamba Pre-training
VI-B1 Random Masking
@@ VI-B2 Masked Pre-training
$\mathbf{X}_{\text{enc}}^{\text{out}}$ $=\mathbf{MLP}(\mathbf{Encoder}(\mathbf{X}_{0}^{\text{vis}}))\in\mathbb{R}^{\mathtt{L}_{\text{vis}}\times\mathtt{D}_{\text{dec}}}$ (5)
$\mathbf{X}_{\text{dec}}^{\text{in}}$ $=\mathbf{Unshuffle}(\mathbf{Concat}(\mathbf{X}_{\text{enc}}^{\text{out}},\mathbf{X}_{\text{mask}}))+\mathbf{E}_{\text{dec}}^{\text{pos}}$
$\mathbf{X}_{\text{dec}}^{\text{out}}$ $=\mathbf{Decoder}(\mathbf{X}_{\text{dec}}^{\text{in}})$
where the $\mathbf{Unshuffle}$ operation restores the original sequence order, and $\mathbf{E}_{\text{dec}}^{\text{pos}}\in\mathbb{R}^{\mathtt{L}\times\mathtt{D}_{\text{dec}}}$ represents decoder-specific positional embeddings. Subsequently, the mean square error (MSE) loss for self-supervised reconstruction is calculated as shown below:
@@
$\mathbf{y}_{\text{rec}}$ $=\mathbf{Shuffle}(\mathbf{X}_{\text{dec}}^{\text{out}})[\mathtt{L}_{\text{vis}}+1:\mathtt{L},\ :\ ]$
@@
where $\mathbf{y}_{\text{real}}$ represents the ground-truth mask tokens, and $\mathbf{y}_{\text{rec}}$ signifies the predicted ones.
VI-C NetMamba Fine-tuning
@@
$\mathbf{X}$ $=\mathbf{Encoder}(\mathbf{X}_{0})\in\mathbb{R}^{\mathtt{L}\times\mathtt{D}_{\text{enc}}}$ (7)
$\hat{\mathbf{y}}$ $=\mathbf{MLP(\mathbf{Norm}(\mathbf{X}[\mathtt{L},\ :]))}$
@@
Takeaway. The unidirectional Mamba architecture is well-suited for processing sequential network traffic data. To acquire generic network domain knowledge, NetMamba is pre-trained by reconstructing masked strides. For adaptation to specific downstream tasks, NetMamba is fine-tuned by minimizing prediction loss.
@@ VII-A1 Datasets
To assess the effectiveness and generalization abilities of NetMamba, we conducted experiments using six publicly available real-world traffic datasets encompassing three main classification tasks.
@@
    Encrypted Application Classification: This task aims to classify application traffic under various encryption protocols. Specifically, the CrossPlatform (Android) [31] and CrossPlatform (iOS) [31] contain 254 and 253 applications respectively. Additionally, we use Tor traffic data from 8 communication categories in ISCXTor2016 [32] and VPN traffic data from 7 communication categories in ISCXVPN2016 [33].
@@
    Attack Traffic Classification: This task aims to identify potential attack traffic, such as Denial of Service (DoS) attacks and brute force attacks. We construct 6 data categories using CICIoT2022[34].
@@
    Malware Traffic Classification: This task aims to distinguish between traffic generated by malware and benign traffic. We use all 20 data categories from the USTC-TFC2016 dataset [35].
@@ VII-A2 Comparison Methods
  4. 4.
    Transformer variants within the NetMamba backbone, including NT-Vanilla and NT-Linear. The former replaces Mamba blocks in NetMamba with vanilla Transformer blocks [36] featuring quadratic complexity, while the latter adopts Linear Transformer blocks [37] with linear complexity.
@@ VII-A3 Implementation Details
The NetMamba architecture features an encoder composed of 4 Mamba blocks and a decoder composed of 2 Mamba blocks. More hyper-parameter details can be found in Table III.
@@
TABLE III: Hyper-Parameter details of NetMamba
@@ VII-B Overall Evaluation
TABLE IV: Comparison Results on CrossPlatform(Android), CrossPlatform(iOS) and CICIoT2022
@@
TABLE V: Comparison Results on ISCXTor2016, ISCXVPN2016 and USTC-TFC2016
@@
We evaluated the performance of NetMamba in categorizing traffic using six publicly available datasets. As shown in Table IV and Table V, NetMamba demonstrates superior performance compared to all baseline methods across five datasets. However, it falls slightly short in comparison to TFE-GNN on the CICIoT2022 dataset, with a marginal 0.15 percentage point difference in accuracy and f1 score. On average, NetMamba achieves accuracy levels ranging from 0.9869 to 0.9993 and f1 scores between 0.9864 and 0.9993. Notably, NetMamba maintains the fewest parameters among all deep learning methods, underscoring its efficient yet effective capabilities in traffic representation learning.
@@ VII-B1 CrossPlatform (Android)
As indicated in Table IV, NetMamba demonstrates significant improvements over existing methods on the CrossPlatform(Android) dataset. Specifically, our model achieves a notable improvement of 4.83% in accuracy and 4.64% in f1 score compared to the state-of-the-art method (ET-BERT). ET-BERT focuses solely on learning traffic representations from packet payloads, overlooking the valuable content information carried by packet headers. In contrast, NetMamba effectively models both header and payload characteristics, leading to a more comprehensive analysis of traffic patterns.
@@ VII-B4 ISCXTor2016, ISCXVPN2016 & USTC-TFC2016
As depicted in Table VI, NetMamba surpasses all existing methods across the three encrypted datasets. Particularly noteworthy is the significantly unstable performance exhibited by methods other than YaTC and YaTC(OF) across different datasets. Given the similarities between NetMamba and YaTC, we contend that a robust traffic representation scheme, which incorporates both header and payload data, alongside a well-designed pre-training task, plays a crucial role in enhancing encrypted traffic analysis capabilities.
@@ VII-C Inference Efficiency Evaluation
Figure 3: The Inference Speed and GPU Memory Comparison
Figure 4: The Inference Efficiency Comparison on Fine-tuning Batch Size
To evaluate the inference efficiency of NetMamba, we conducted experiments comparing its speed and GPU memory consumption with existing deep learning methods. Speed is measured as the number of traffic data samples processed by the model per second: packets for ET-BERT and flows for the others. As shown in Figure 4(a), NetMamba achieves the highest inference speed across various input batch sizes, with improvements ranging from 1.22 to 60.11 times. This advantage is particularly notable due to the substantial model parameters and inefficient model architecture design present in models such as ET-BERT, TFE-GNN, and FS-Net. Even when compared with models possessing similar parameter counts, NetMamba continues to outperform NT-Vanilla, YaTC and its faster variant. This superiority is primarily attributed to Mamba’s lower computational complexity compared to Transformer models. Given a token sequence $\mathbf{X}\in\mathbb{R}^{1\times\mathtt{L}\times\mathtt{D}}$ and the default setting $\mathtt{E}=2\mathtt{D},\ \mathtt{N}=16$, the computational complexities of vanilla or linear attention in Transformers and SSM in Mamba are as follows:
$\Omega(\text{Vanilla-Attention})=4\mathtt{L}\mathtt{D}^{2}+2\mathtt{L}^{2}\mathtt{D}$ (9)
$\Omega(\text{Linear-Attention})=3\mathtt{L}\mathtt{D}^{2}+2\mathtt{L}\mathtt{D}$ (10)
@@
Self-attention exhibits quadratic complexity to the sequence length $\mathtt{L}$, whereas SSM operates linearly. This computational efficiency makes NetMamba more scalable than Transformer-based models like YaTC and ET-BERT. Although both NT-Linear and NetMamba achieve linear complexity and have similar parameter counts, based on Equation 10 and (11), the SSM requires less computational cost when $\mathtt{D}>42$. Given our use of $\mathtt{D}=256$, NT-Linear’s slower inference speed compared to NetMamba is reasonable.
In Figure 4(b), NetMamba demonstrates lower GPU memory consumption than most models, except FS-Net, YaTC and NT-Linear, when using large batch sizes. FS-Net’s reliance on RNNs, which require linear memory relative to sequence length, reduces memory costs but results in slower inference and poorer classification performance. YaTC reduces memory usage by shortening input sequence length through a model forward trick. Without such an optimization, YaTC(OF) consumes up to four times more GPU memory than NetMamba As shown in the subsequent ablation study, NT-Linear exhibits unstable classification performance due to over-compression of standard attention mechanisms. Compared to other baselines, NetMamba achieves improved memory efficiency primarily by customizing GPU operators that minimize the storage of extensive intermediate states and conduct recomputation during the backward pass.
When the input batch size is set to 64 (the value used in fine-tuning), as depicted in Figure 4, NetMamba exhibits an improvement in speed, being 2.24 times faster than the best baseline, YaTC(OF). Apart from FS-Net, memory-optimized YaTC, and NT-Linear, NetMamba surpasses other methods in terms of GPU memory utilization. In summary, NetMamba achieves the highest inference speeds among all deep learning methods while maintaining comparably low memory usage.
TABLE VI: Ablation Study of NetMamba on All Datasets
@@
w/ Bidirectional Mamba1 0.9852 0.9851 0.9793 0.9793 0.9974 0.9974 0.9986 0.9986 0.9935 0.9935 0.9987 0.9987
w/ Cascading Mamba1 0.8939 0.8942 0.9470 0.9455 0.9826 0.9826 0.9952 0.9952 0.9646 0.9643 0.9938 0.9937
w/ Vanilla Transformer2 0.9815 0.9815 0.9755 0.9755 0.9979 0.9980 0.9986 0.9986 0.9928 0.9927 0.9997 0.9997
w/ Linear Transformer2 0.9806 0.9820 0.9871 0.9871 0.8954 0.8998 0.9986 0.9986 0.8295 0.8398 0.4014 0.4226
@@
    Substituted unidirectional Mamba blocks with either bidirectional blocks [12] or cascading ones[30].
@@
    Replaced Mamba blocks with either vanilla [36] or linear[37] Transformer blocks, termed NT-Vanilla and NT-Linear, respectively.
  • 3
    Changed the 1-dimensional stride cutting to 2-dimensional patch splitting.
Figure 5: The Performance Comparison on Few-Shot Settings